OpenStudy (anonymous):

A sporting goods store offers a 40% discount on all golf clubs. Rocco spent 20% of the money in his savings account on a golf putter. He paid $48 for the putter after the discount. How much money did Rocco have in his savings account before buying the putter?

OpenStudy (nincompoop):

it says that he spent 20% of his money on a golf putter, and then it continued to say that the amount was $48. so what you need to understand and ask: $48 is 20% of what amount ?

OpenStudy (nincompoop):

would it be logically consistent to formulate: \[\frac{ 48 }{ x }=\frac{ 20 }{ 100 }\] 48 is .20 of amount x

OpenStudy (nincompoop):

then all we have to do is solve for x people suggest the "cross multiply" solution, which is okay.

OpenStudy (nincompoop):

well if you are not going to respond and just sit there for us to solve your problem, you might as well drop out of school now.

OpenStudy (abb0t):

Mayb we should call the ambulance..

OpenStudy (anonymous):

I'm not quite understanding the way of how the cross multiplication works with this

OpenStudy (nincompoop):

okay let us skip cross multiplication. in that case, how would you solve for x?

OpenStudy (nincompoop):

when dealing with a variable to solve that is located in the numerator, we can always take the reciprocal of both the equation so it would be easier to look at this this way: \[\frac{ x }{ 48 } = \frac{ 100 }{ 20 }\]

OpenStudy (anonymous):

Okay so it means that we have to put a number on top of the 48 that would make the fraction equivalent to \[100/20\] ?

OpenStudy (anonymous):

YOU GUYS SEEM TO FORGET THE 40% DISCOUNT

OpenStudy (nincompoop):

we are solving for x and we have x/48 @Samour it does not matter how much was the putter's original price before the discount. the central question here is "How much money did Rocco have in his savings account before buying the putter?"

OpenStudy (nincompoop):

the putter's price is marked at 40% discount. it will be sold in that discounted amount no matter what and anyone who will buy it will pay %48 bucks. the original price is a smoke screen, it is not relevant to what we want to solve.